INAF LBC+MODS queue observing and AO PISCES, 2012 June 21-22

Observer: C. Arcidiacono, R. Briguglio, K. Boutsia, M. Faccini, F. Vitali
Telescope Operator: S. Allanson
Support Astronomer: D. Thompson
Instrument: LBC + PISCES

Summary

Clear skies at the beginning of the night and the average wind is 6m/s.
The night is dedicated to Pisces. We open up early (1:30UT) to do narrow band flats.
The seeing and the wind have been decent for the whole night.
No major problems with the AO system. Once the msgd crashed and they had to restart the secondary. We also had a couple of "shell rips" but we did not loose a lot of time.
Three more projects were completed tonight with Pisces.
At the end of the night, during twilight, we switched to LUCI to get some calibration frames in J and Ks of M15, for an older AO project.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Fiore Mrk231 J, Ks 20min+sky / 2min Ks COMPLETED
Barbieri KIC10864656 J,H,Ks ~4min x fliter + sky COMPLETED
Ragazzoni PCygni FeII, 2.14um, Ks ~30min x filter/ 5min Ks + sky COMPLETED
Bono M15 J,Ks 3min x filter calibrations with LUCI for an AO project

Details

01:30 2.14continuum flats with Pisces

03:15 Telescope is collimater and we can send the first preset (Fiore - Mrk231)
  • dimm seeing 1.1" - GCS seeing 1.1"
  • loop closed

03:26 start with science block
  • failed during offset - have to close the loop manually

03:29 send the preset again
  • the quality of the first images was not very good
  • we had to manually adjust the autogain

03:40 repeat block of images

03:43 no star on the guider for an offset position - but the loop is closed so we keep taking data
  • dimm seeing 0.9"

04:04 the loop opened but we are taking sky frames

04:21 once we offset back tothe target, it finds the star on the pyramid, the loop closes and we can go on

05:01 ACE preset to go back to the target after the sky frames

05:04 not Wavefrontsensing - power cycle the AZcam server again??

05:15 send preset again with autocollimate - now it seems to be working

05:35 we obtain also a block of Ks images because the seeing is better than the previous nights.

05:50 preset cancelled because the shell ripped!!

05:57 send active preset for calibration star AO534

06:17 Slewing to Barberi (KIC1086)
  • preset cancelled because the shell ripped
  • MessageDemon off-line

06:32 problem solved - sending a preset

06:43 science block starts
  • we take also some longer exposures (10sec) where the center is slightly saturated

07:24 Slew to PCygni (Ragazzoni)
  • test images to decide the dither pattern

07:59 we start the science block

08:18 loop open - try to close it manually

08:28 loop opened

08:31 loop closed again - we start the script again with less offsets

08:45 send the offset with 30deg rotations
  • the loop opened

08:47 wrong PA in the sent offset
  • repeat preset again

08:57 start science block at 30deg

09:39 acquireRef failed while we were sending a preset in track mode

10:03 sky frames for the PCygni project (Ragazzoni)

10:14 SWITCH TO LUCI

10:20 Slew to M15 for photometric calibration frames with LUCI

10:28 start script for J band - 10 sec exposures luci.20120622.0003.fits-0015.fits
  • dimm seeing 1.0" - GCS seeing 0.6"

10:45 start script for Ks band - 5sec exposures luci.0016-0028
  • images on Luci are elongated

11:00 start script for J band - 5sec exposures luci.0029-0041

11:13 start script for Ks band - sky frames a bit offset luci.0042-0054

11:28 start Ks flats (low - 2000counts) luci.0055-0064

11:32 J flats (low -1000counts/high -5000counts) luci.0065-0084

11:55 start Ks flats (high - 5000counts) luci.0085-0094

11:58 park telescope at zenith to get the Pisces flats ( J,H,Ks)

12:25 pisces stopped while we were taking flats - telnet connection is not responding

12:55 send scripts for darks (both for LUCI and Pisces)

13:26 END OF RUN!!

-- KonstantinaBoutsia - 22 Jun 2012
Topic revision: r6 - 12 Jul 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback